OpenStudy (lifeisadangerousgame):

How many solutions are there to the following system of equations? 5x + 3y = 7 2x – y = 2

OpenStudy (anonymous):

two solutions

OpenStudy (callisto):

For a equation in this form; Ax + By + C =0 slope of the line = -A/B y-int = -C/B 5x + 3y = 7 => 5x+3y-7 = 0 -(1) 2x – y = 2 => 2x-y-2 =0 -(2) So, slope of equation (1) = -5/3 y-int of equation (1) = -(-7/3) = 7/3 and slope and y-int of equation (2) = ...??? If they have same slope, different y-int, they are // lines => no intercepts If they have same slope, and same y-int, they are the same line => infinitely many intercepts If they have different slope, they are not // lines => 1 intercept Can you work out the answer now?

OpenStudy (callisto):

Of course, another way to do it is to solve the equations :|

OpenStudy (unklerhaukus):

how can their be two solutions to points on different straight lines?

OpenStudy (unklerhaukus):

clearly m is different , hence there is only one solution

OpenStudy (anonymous):

the given lines are having two solutions as they aquire diffrent slope..:) so their values of x and y are different..:)

OpenStudy (anonymous):

two solutions? that would be a miracle

OpenStudy (anonymous):

5x+3y=7 2x- y =2 take the coeficiets of first terms multiply first with secod and second with first alternatively you will get 10x+6y=14 10x-5y=10 change the sign 10x+6y=14 -10x+5y=-10 ========= 0+11y=4 y=4/11 now put this value in any equation to get value of x

OpenStudy (unklerhaukus):

can you draw two straight lines with different slops that intersect twice @sugargurl ?

OpenStudy (anonymous):

the choices are 1) one solution (lines intersect at a point) 2) no solutions (lines are parallel, and do not intersect) 3) infinitely many solutions (it is really the same line) two solutions is not an option

OpenStudy (anonymous):

yes unklerhaukas u r correct sorry i was wrong thanks for correcting me,,:)

OpenStudy (anonymous):

The first one answer is correct

OpenStudy (anonymous):

Graph of your question

OpenStudy (anonymous):

d=(v^2/g)*sin2theta

OpenStudy (lifeisadangerousgame):

Well the answers were: A) 0, B) 1, C) 2, D) no solutions it's kinda late but yea thanks everyone :)